OpenStudy (anonymous):

Find the slope-intercept form for the line satisfying the following conditions. x-intercept=2, y intercept =4/5

OpenStudy (anonymous):

Why is it negative ?

OpenStudy (anonymous):

Shouldn't it just be y= 4/15+ 4/5

OpenStudy (anonymous):

the line would be \(\frac{x}{2}+\frac{y}{\frac{4}{5}} =1\) rearrange to point intercept form. it is neg. gradient

OpenStudy (anonymous):

Ok so what if its x intercept is 3 and y is 1/3

OpenStudy (anonymous):

the line would be \(\frac{x}{3} +\frac{y}{\frac{1}{3}} =1\) rearrange again. the line has a neg gradient. I believe you've noticed the correlation?

OpenStudy (anonymous):

So y= -1/9+1/3

OpenStudy (anonymous):

y= -x/9+1/3 yup. u've got it.

OpenStudy (anonymous):

Cool thanx

OpenStudy (anonymous):

you're welcome :)